OpenStudy (anonymous):

Does the set of numbers 13, 21, and 24 form a Pythagorean triple? Explain. (1 point)

OpenStudy (anonymous):

yes; 132 + 212 ≠ 242 no; 132 + 212 = 242 no; 132 + 212 ≠ 242 yes; 132 + 212 = 242

OpenStudy (anonymous):

help please?

OpenStudy (raden):

just prove that 24^2 = 13^2 + 21^2 if not equal, it cant form a Pythagorean triple

OpenStudy (anonymous):

Idk how to do it?

OpenStudy (raden):

hint : 24^2 = 24 * 24 = 576

OpenStudy (anonymous):

so it'd be no right?

OpenStudy (raden):

how do u know that ?

OpenStudy (raden):

24^2 = 24 * 24 = 576 21^2 = 21 * 21 = 441 13^2 = 13 * 13 = 169

OpenStudy (raden):

now, is it true 576 = 441+169 ?

OpenStudy (anonymous):

so which answer would I choose out of my choices?

OpenStudy (anonymous):

it's not right.

OpenStudy (raden):

576 = 441+169 576 = 610 (it is not right) i576 ≠ 610 yes, u are right

OpenStudy (raden):

miss for i :)

OpenStudy (anonymous):

which answer then? a b c or d? yes; 132 + 212 ≠ 242 no; 132 + 212 = 242 no; 132 + 212 ≠ 242 yes; 132 + 212 = 242

OpenStudy (raden):

C

OpenStudy (raden):

any else ?